OpenStudy (anonymous):

The speed of sound, V, near Earth's surface can be represented by this equation, V=20sqt(t+273) where t is the temperature in degrees Celsius. Find the temperature when V = 380 meters per second. Round to the nearest whole number.

OpenStudy (anonymous):

88

OpenStudy (anonymous):

\[V=20\sqrt(t+273)\] \[380=20\sqrt(t+273)\] \[19=\sqrt(t+273)\] \[19^{2}=361\] Square both sides and get \[361=273+t\] \[361-273=t\] \[88=t\]